X Company currently makes a part and is considering buying it next year from a company that has offered to supply it for $17.56 per unit. This year, total costs to produce 58,000 units were:

Direct materials $377,000
Direct labor 301,600
Variable overhead 237,800
Fixed overhead 69,600


If X Company buys the part, it can avoid $24,360 of the fixed overhead. The resources that will become idle if they choose to buy the part can be used to increase production of another product, resulting in additional total contribution margin of $70,000.

The marketing manager is uncertain what demand will be next year. What level of demand will make the company indifferent between making the part and buying it?
Answer #1

Answer is 53614 units

Make
Direct Material 377000
Direct Labor 301600
Variable Overhead Cost 237800
Total Variable Cost 916400
No of Units Produced 58000
Variable Cost per unit 15.80
At indifferent cost under to options are Equal
$70000 Opportunity cost
Let X no of units ie Indifferent Point
15.80X+69600+70000-=17.56X+(69600-24360)
Solving the above equation
15.8x+139600=17.56x+45240
X=94360/1.76 =53614 units